OpenStudy (anonymous):

my answer: y = 5/2x + 5 Change the given equation into slope-intercept form. 2x + 5y = -10

OpenStudy (compassionate):

2x + 5y = -10 5y = 2x - 10 y = 5/2x - 10 ;)

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

thats not an option

OpenStudy (compassionate):

Oops, made an error ~ 2x + 5x = -10 5x = -2x - 10 y= -(5/2x) - 5

OpenStudy (anonymous):

you made another error

OpenStudy (compassionate):

Dang t. It's like mistake monday. y = -2x/5 - 5! THERE! I'm 100% that's correct.

OpenStudy (anonymous):

-5/2x + 5?

OpenStudy (compassionate):

DANG IT y = -2x/5 - 2 !!!!!!!!

OpenStudy (compassionate):

Don't laugh at me. My brain is having a retard moment ;-;

OpenStudy (anonymous):

hahahahaahahahahhahahah :))))

OpenStudy (compassionate):

But I'm sure my last post was correct!

OpenStudy (anonymous):

there you go sweetie

OpenStudy (anonymous):

can you help me plz Compassionate you did not finish